OpenStudy (anonymous):

Given the equation Square root of 2x plus 1 = 3, solve for x and identify if it is an extraneous solution. It would start with - 3 correct?

OpenStudy (anonymous):

I'm after an explanation rather than answer here.

OpenStudy (anonymous):

√2x+1=3, right?

OpenStudy (anonymous):

Yes, @AlexDg3

OpenStudy (welshfella):

does the square root apply to the 2x only or is it sqrt(2x + 1)?

OpenStudy (anonymous):

\[\sqrt{2x+1} = 3\]

OpenStudy (anonymous):

The first thing you have to do is square both sides of the equation to get 2x+1=9

OpenStudy (anonymous):

How would 2x not be affected by the SQ?

OpenStudy (anonymous):

because its under the square root

OpenStudy (anonymous):

When you square it, it cancels the square root

OpenStudy (anonymous):

Ah. So anything that isnt under a SQR is affected?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah. So anything that isnt under a SQR is affected?

OpenStudy (anonymous):

after you square it and get 2x+1=9, you solve for x

OpenStudy (anonymous):

Ok, well this is looking for any extraneous answers, which in what I've understood in layman's terms is an answer that works on one half but not another.

OpenStudy (anonymous):

x = 4, solution is extraneous x = 4, solution is not extraneous x = 5, solution is extraneous x = 5, solution is not extraneous thats what I was given.

OpenStudy (anonymous):

And it would be x=5 not extraneous correct?

OpenStudy (anonymous):

I don't think so. There is only one x, and it can only equal 4.

OpenStudy (anonymous):

I think.

OpenStudy (anonymous):

But isn't it square root then sum up what you have? (being after you square it and get 2x+1=9)

OpenStudy (anonymous):

No, its four. you're right. I did it wrong.

OpenStudy (welshfella):

4 not extraneous because x = 4 satisfies the equation

OpenStudy (anonymous):

I had already figured it was non extraneous.